OpenStudy (anonymous):

How do I simplify the radical expression 6+square root 3 over 5-square root 3?

OpenStudy (anonymous):

\[\frac{6+\sqrt3}{5-\sqrt3}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i think "simplify" in this case actually means "rationalize the denominator" because there is no such mathematical operation as "simplify" multiply top and bottom by the conjugate of the denominator the conjugate of \(5-\sqrt3\) is \(5+\sqrt3\) and this works because \[(5-\sqrt3)(5+\sqrt3)=5^2-3=22\]

OpenStudy (anonymous):

you get \[\frac{6+\sqrt3}{5-\sqrt3}\times \frac{5+\sqrt3}{5-\sqrt3}\]\[=\frac{(6+\sqrt3)(5+\sqrt3)}{22}\]

OpenStudy (anonymous):

then you have to multiply out in the numerator

OpenStudy (anonymous):

you got if from here?

OpenStudy (anonymous):

I think so. the answer that I got is \[\frac{ 1\sqrt{3} }{ 2 }\]